OpenStudy (anonymous):

INTEGRATE---- pLEASE CHECK FOR ME WHETHER MY ANSWER IS RIGHT OR WRONG (integral) (dx/(x(x^2+1)^2)) i get lnx +1/2x-1/(x^2+1)

OpenStudy (anonymous):

my answer \[lnx+\frac{ 1 }{ 2x }-\frac{ 1 }{ x^2+1 }\]

OpenStudy (anonymous):

maybe you still not clear with the q \[\int\limits{}^{}\frac{ dx }{x(x^2+1)^2}\]